# Prague, Czech

### v3

@@ 4.1 Mixed-Precision GMRES-based Iterative Refinement
We instantiate the general RL framework for the GMRES-IR method (doi:10.1137/17M1140819, ), selecting precisions for four computational steps to solve linear systems $Ax=b$, reconciling the competing demands of numerical accuracy and computational efficiency. Consider a linear system $Ax=b$, where $A\in\mathbb{R}^{n\times n}$ is non-singular, and $b\in\mathbb{R}^{n}$. GMRES-IR iteratively refines an initial solution $x_{0}$ as:
@@
We note that it is also possible to use mixed precision within the GMRES method itself; see, e.g., vi22 . For simplicity, we restrict ourselves here to GMRES in uniform precision. The goal is to use RL to select an action $\mathbf{a}=(u_{f},u_{r},u_{g},u)$ to minimize the final residual norm $\|b-Ax_{\text{solve}}\|_{2}$ and the total number of iterations $T$, subject to a trade-off between $T_{\text{max}}$ and the convergence tolerance $\tau$.
Inspired by stopping criteria from 10.1145/1141885.1141894 , we consider the algorithm converged if:
@@ 4.2 Contextual Bandit Formulation for Precision selection
where $\kappa(A)$ is the condition number of $A$ that can be approximated via an efficient algorithm (e.g., Hager-Higham Hager:1984:CE ; Higham:1987:SCN ), $\delta_{c}>0$ ensures numerical stability in logarithmic scaling, $\|A\|_{\infty}=\max_{i}\sum_{j}|a_{ij}|$ is the infinity norm, and $\delta_{n}>0$ prevents numerical issues. These features characterize the numerical stability and scaling of the linear system and are fast to compute.
@@
The action $\mathbf{a}\in\mathcal{A}$ is a tuple $\mathbf{a}=(u_{f},u,u_{g},u_{r}),\quad u_{i}\in\mathcal{A}_{1}$ where $\mathcal{A}_{1}=\{u_{1},\ldots,u_{m}\}$ is the set of precisions, and $m=|\mathcal{A}_{1}|$. Following (9), the action space $\mathcal{A}=\mathcal{A}_{1}^{4}$ is constrained by:
$u_{f}\leq u\leq u_{g}\leq u_{r},$ (18)
where $\leq$ denotes ordering by increasing significand bits.
The reward $r(s,\mathbf{a})$ is designed to lead RL to choose precision configurations that lead the linear solvers to have an acceptable level of converged iterations as well as accuracy, it has the form:
@@
    $f_{\text{accuracy}}(s_{d},\mathbf{a})$ penalizes the error in the solution relative to the true solution and the normalized residual. It is computed as:
    $f_{\text{accuracy}}(s_{d},\mathbf{a})=-\log_{10}(\max(\text{err}_{\text{distance}},\mathbf{C}_{1}))-\log_{10}(\max(\text{err}_{\text{normalized}},\mathbf{C}_{1})).$ (22)
    If $\text{err}_{\text{distance}}>1.0\text{ or }\text{err}_{\text{normalized}}>1.0$ we set $f_{\text{accuracy}}(s_{d},\mathbf{a})=\mathbf{C}_{2}$. Empirically, we choose $\mathbf{C}_{1}=10^{-10}$ and $\mathbf{C}_{2}=5.0$.
  • Penalty: We use $f_{\text{penalty}}$ term to denote the penalty arising from computational cost, e.g., iterations, or failure steps such as LU factorization or stagnation calculation. In our empirical study, it is sufficient to use the converged iterations to model the penalty term; that is, the more iterations we use, the larger the penalty term will be. In this context, we can either use the total number of iterations for iterative refinement or the total (inner-solve) GMRES iterations used for iterative refinement, denoted by $T_{\text{iter}}$. The following definition is used in our context:
@@
1:Input: Training systems $\{(A_{j},b_{j})\}_{j=1}^{N}$, GMRES-IR method $\mathcal{M}$, precision set $\mathcal{A}_{1}=\{u_{1},\ldots,u_{m}\}$, reward function $r$, episodes $T$, tolerance $\tau$, max iterations $T_{\text{max}}$, learning rate $\alpha$, discount factor $\gamma$, minimum exploration probability $\epsilon_{\text{min}}$, bins $\{\Phi_{1},\Phi_{2}\}$, top-$k$ parameter $k_{\text{top}}$
2:Compute $\mathcal{A}_{\text{reduced}}$ in terms of $\mathcal{A}_{1}$
3:Initialize Q-table $Q(s_{d},\mathbf{a})$ for $s_{d}\in\mathcal{S}_{d}$, $\mathbf{a}\in\mathcal{A}_{\text{reduced}}$
@@
20:   Compute next system index $j^{\prime}=(j\bmod N)+1$, context $s^{\prime}$, and discretized state $s_{d}^{\prime}$
21:   Update Q-table:
@@
23:end for
24:Inference: For new system $(A,b)$, compute $s=[\log_{10}(\max(\kappa(A),\delta_{c})),\log_{10}(\max(\|A\|_{\infty},\delta_{n}))]$, discretize $s_{d}$, select $\mathbf{a}=\arg\max_{\mathbf{a}\in\mathcal{A}_{\text{reduced}}}Q(s_{d},\mathbf{a})$, solve with $\mathcal{M}$ using $\mathbf{a}$
@@ 5 Experiments
We conducted experiments to evaluate the proposed RL framework for precision selection in GMRES-IR for solving linear systems $p=(A,b)\in\mathcal{P}$, where $A\in\mathbb{R}^{n\times n}$ is a non-singular matrix and $b\in\mathbb{R}^{n}$. The systems are derived from random generated matrices. Our code is simulated in Python and uses Pychop (carson2025, ) for precision emulation.
We generated problems $Ax=b$ using two settings, each designed to evaluate the RL-driven precision selection for the GMRES-IR solver under different numerical conditions versus a full double precision routine (in the following, we use FP64 and double precision interchangeably). The agent with learning rate $\alpha$ of $0.5$ is trained within a custom Gym environment (brockman2016openai, ) with 100 episodes, which simulates the process of solving a linear system using the GMRES-IR solver. To reduce the action space and accelerate computation, we adopted the reduction strategy outlined in Section 3.3 and further pruned it by selecting one-fourth of the valid precision combinations from the episode number combinations.
@@ 5.1 Problem Setup
where $|\mathcal{R}_{j}|$ is the cardinality $\mathcal{R}_{j}$, and the threshold $\tau_{j}$ is scaled by the condition number to account for problem difficulty, aligning with the error bounds in 10.1145/1141885.1141894 .
@@ 5.2 Simulations on Dense Systems
Table 3: Average Floating-point Types Selected Across Condition Ranges for Dense Systems.
Weight Setting Condition Range Num BF16 TF32 FP32 FP64
$\tau=10^{-6}$
RL ($W_{1}$) $10^{0}\!-\!10^{1}$ 33 $0.00$ $0.00$ $0.33$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.00$ $0.00$ $0.43$ $3.57$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.00$ $0.86$ $3.14$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.33$ $3.67$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.78$ $3.22$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
RL ($W_{2}$) $10^{0}\!-\!10^{1}$ 33 $0.33$ $0.00$ $0.00$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.19$ $0.76$ $0.10$ $2.95$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.00$ $0.86$ $3.14$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.23$ $0.23$ $3.54$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.33$ $3.67$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.78$ $3.22$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
$\tau=10^{-8}$
RL ($W_{1}$) $10^{0}\!-\!10^{1}$ 33 $0.00$ $0.00$ $0.33$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.00$ $0.00$ $0.43$ $3.57$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.29$ $0.29$ $3.43$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.20$ $3.80$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.11$ $3.89$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
RL ($W_{2}$) $10^{0}\!-\!10^{1}$ 33 $0.33$ $0.00$ $0.00$ $3.67$
$10^{1}\!-\!10^{2}$ 2121 $0.14$ $0.76$ $0.10$ $3.00$
$10^{2}\!-\!10^{3}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{3}\!-\!10^{4}$ 77 $0.00$ $0.29$ $0.29$ $3.43$
$10^{4}\!-\!10^{5}$ $13$ $0.00$ $0.00$ $0.23$ $3.77$
$10^{5}\!-\!10^{6}$ $15$ $0.00$ $0.00$ $0.33$ $3.67$
$10^{6}\!-\!10^{7}$ 99 $0.00$ $0.00$ $0.78$ $3.22$
$10^{7}\!-\!10^{8}$ 99 $0.00$ $0.00$ $0.22$ $3.78$
$^{8}\!-\!10^{9} 1010$ $0.00$ $0.00$ $0.00$ $4.00$
Table 5.2 reports error metrics and iteration counts for the linear solve, organized by condition number ranges: low ($10^{0}$ to $10^{3}$), medium ($10^{3}$ to $10^{6}$), and high ($10^{6}$ to $10^{9}$). Table 3 details the precision types selected within finer condition number intervals.
@@
In terms of convergence, the conservative $W_{1}$ policy achieves universal convergence with $\xi=100\%$ across all settings. By comparison, the aggressive $W_{2}$ policy maintains $100\%$ success at $\tau=10^{-6}$ but shows reduced robustness at the stricter tolerance $\tau=10^{-8}$, with $\xi$ decreasing to $89.2\%$ in the low-$\kappa$ tests (RL fails in a few test samples), likely because the RL had insufficient exposure to training data in this specific regime. With respect to accuracy, $W_{1}$ closely matches the FP64 baseline, producing forward error with the same order of magnitude and comparable backward errors. Precision lower than FP32 is never used in $W_{1}$. In contrast, $W_{2}$ exhibits a clear trade-off: BF16 and TF32 start to be used within low-to-medium $\kappa$ ranges, where smaller condition numbers theoretically permit lower-precision computations with decreased error amplification. In $W_{2}$, it permits the use of low precisions with substantially higher errors (e.g., ferr $\approx 10^{-7}$) to optimize the reward function, as long as the convergence quality is guaranteed. This behavioral distinction manifests in computational effort as well. The average number of outer iterations remains stable ($\approx 2.0$ to $2.35$) for both policies. However, the inner solver cost varies. The $W_{2}$ policy requires slightly more inner GMRES iterations (e.g., $8.00$ compared to $2.70$ for $W_{1}$ in low-$\kappa$ bins). This increase is needed to offset the growing rounding errors introduced by lower-precision formats. The visualization of ferr and total GMRES iterations used for each test sample in $W_{2}$ is shown in Figure 2. The average performance degradation in Table 5.2 from the $W_{2}$ is offset by a few instances that deviate from the identity line due to the aggressive strategy with reduced-precision preference.
The precision selection patterns in Table 3 clarify these performance outcomes. In the low-$\kappa$ regime ($10^{0}$ to $10^{2}$), $W_{1}$ employs a conservative combination of FP64 and FP32, entirely avoiding lower precisions. In contrast, $W_{2}$ leverages BF16 and TF32 (with usage frequencies of approximately $0.33$ and $0.76$, respectively), reflecting a focus on cost-efficiency. As the condition number increases, both strategies converge; beyond $10^{6}$, both policies adopt an FP64-dominant strategy and avoid the use of lower precisions.
@@ 5.3 Simulations on Sparse Systems
In this simulation, we emulated sparse linear systems following the experiment described in hausner2024neural . A matrix $A_{0}\in\mathbb{R}^{n\times n}$ is constructed with $\text{nnz}(A_{0})=\lfloor\lambda_{s}n^{2}\rfloor$ non-zero entries, where $\lambda_{s}>0$ is the sparsity parameter. The entries are sampled from a standard normal distribution:
@@
The evaluation of sparse linear systems presents a distinct contrast to the dense case, primarily driven by the inherent numerical difficulty posed by the very ill-conditioned dataset. As depicted in Table 4, the test matrices exhibit high condition numbers between $1.02{\times}10^{8}$ and $7.86{\times}10^{9}$, which places linear solve in a regime where numerical stability is precarious, leaving limited room for precision reduction. This test effectively demonstrates whether the RL agent can effectively avoid numerical instabilities in its precision selection.
@@
5.4 Ablation Study
To further verify the effectiveness of our RL modeling for precision selection. We focus on the converged total GMRES iterations taken for the iterative refinement and conduct an ablation study on training with (19) without $f_{\text{penalty}}$—the penalty on iteration terms.
The results are presented in Table 5.4, together with Figure 3 showing precision types chosen by RL. Similarly to above, we observed that between low- and medium-conditioned linear systems, RL strategically selects aggressive reduced-precision steps, whereas in high-conditioned linear systems it adopts a more conservative strategy for precision selection. The removal of it $f_{\text{penalty}}$ results in increasing iterations for the RL agent to achieve comparable accuracy, which means the iterative refinement uses extra iterations to compensate for the accuracy loss resulting from using increased low-precision steps. The results indicate the importance of the reward function modeling for RL training and demonstrate the effectiveness of RL for precision selection under proper modeling.
Table 7: Average Performance Metrics Across Condition Ranges for Dense Systems (Ablation study).

### v4

@@ 4.1 Mixed-Precision GMRES-based Iterative Refinement
We instantiate the general RL framework for the GMRES-IR method (Carson and Higham, 2018), selecting precisions for four computational steps to solve linear systems $Ax=b$, reconciling the competing demands of numerical accuracy and computational efficiency. Consider a linear system $Ax=b$, where $A\in\mathbb{R}^{n\times n}$ is non-singular, and $b\in\mathbb{R}^{n}$. GMRES-IR iteratively refines an initial solution $x_{0}$ as:
@@
It is worth noting that mixed precision can also be applied directly within the GMRES algorithm itself (see, for example, Vieublé (2022)). In this work, however, we limit our attention to GMRES implemented with a single, consistent precision for simplicity.
The goal is to use RL to select an action $\mathbf{a}=(u_{f},u,u_{g},u_{r})$ to minimize the final residual norm $\|b-Ax_{\text{solve}}\|_{2}$ and the total number of iterations $T$, subject to a trade-off between $T_{\text{max}}$ and the convergence tolerance $\tau$.
Inspired by stopping criteria from Demmel et al. (2006), we consider the algorithm converged if:
@@ 4.2 Contextual Bandit Formulation for Precision selection
where $\kappa(A)$ is the condition number of $A$ that can be approximated via an efficient algorithm (e.g., Hager-Higham Hager (1984); Higham (1987) and GNNs Carson and Chen (2026)), $\delta_{c}>0$ ensures numerical stability in logarithmic scaling, $\|A\|_{\infty}=\max_{i}\sum_{j}|a_{ij}|$ is the infinity norm, and $\delta_{n}>0$ prevents numerical issues. These features characterize the numerical stability and scaling of the linear system and are fast to compute.
@@
The action $\mathbf{a}\in\mathcal{A}$ is a tuple $\mathbf{a}=(u_{f},u,u_{g},u_{r}),\quad u_{i}\in\mathcal{A}_{1}$ where $\mathcal{A}_{1}=\{u_{1},\ldots,u_{m}\}$ is the set of precisions, and $m=|\mathcal{A}_{1}|$. Following (12), the action space $\mathcal{A}_{\text{reduced}}\subseteq\mathcal{A}$ is the top $k$ elements of $\mathcal{A}=\mathcal{A}_{1}^{4}$ is constrained by $u_{f}\leq u\leq u_{g}\leq u_{r}$ where $\leq$ denotes ordering by increasing significand bits.
The reward $r(s,\mathbf{a})$ is designed to lead RL to choose precision configurations that lead the linear solvers to have an acceptable accuracy and number of iterations until convergence, it has the form:
@@
    $f_{\text{accuracy}}(s_{d},\mathbf{a})$ penalizes the error in the solution relative to the true solution and the normalized residual. We define the accuracy-related reward as a piecewise function based on the forward and normalized backward errors:
    $f_{\text{accuracy}}(s_{d},\mathbf{a})=-C_{1}\left(\min\left(\log_{10}(\max(\texttt{ferr},\varepsilon)),\,\theta\right)+\min\left(\log_{10}(\max(\texttt{nbe},\varepsilon)),\,\theta\right)\right),$ (24)
    where ferr and nbe are defined in (17), $\varepsilon$ is a small constant (e.g., $10^{-10}$), and $\theta>0$ is a truncation threshold that limits the reward when the errors are large. Often, setting $\theta=2.5$ is applicable to most cases.
  • Penalty: We use the term $f_{\text{penalty}}$ to denote the penalty arising from computational cost, e.g., iterations or failure steps such as LU factorization or stagnation calculation. In our empirical study, it is sufficient to use the converged iterations to model the penalty term; that is, the more iterations we use, the larger the penalty term will be. In this context, we can either use the total number of iterations for iterative refinement or the total (inner-solve) GMRES iterations used for iterative refinement, denoted by $T_{\text{iter}}$. The following definition is used in our context:
@@
1:Input: Training systems $\{(A_{j},b_{j})\}_{j=1}^{N}$, GMRES-IR method $\mathcal{M}$, precision set $\mathcal{A}_{1}=\{u_{1},\ldots,u_{m}\}$, reward function $r$, episodes $T$, tolerance $\tau$, max iterations $T_{\text{max}}$, learning rate $\alpha$, minimum exploration probability $\epsilon_{\text{min}}$, bins $\{\Phi_{1},\Phi_{2}\}$, top-$k$ parameter $k_{\text{top}}$
2:Compute $\mathcal{A}_{\text{reduced}}$ in terms of $\mathcal{A}=\mathcal{A}_{1}^{4}$ with $k_{\text{top}}$ elements satisfying (11)
3:Initialize tabular action-value estimator $Q(s_{d},\mathbf{a})$ for $s_{d}\in\mathcal{S}_{d}$, $\mathbf{a}\in\mathcal{A}_{\text{reduced}}$
@@
20:   Update tabular action-value estimator:
@@
23:Inference: For new system $(A,b)$, compute $s=[\log_{10}(\max(\kappa(A),\delta_{c})),\log_{10}(\max(\|A\|_{\infty},\delta_{n}))]$, discretize $s_{d}$, select $\mathbf{a}=\arg\max_{\mathbf{a}\in\mathcal{A}_{\text{reduced}}}Q(s_{d},\mathbf{a})$, solve with $\mathcal{M}$ using $\mathbf{a}$
@@ 5 Experiments
We conducted experiments to evaluate the proposed RL framework for precision selection in GMRES-IR for solving linear systems $p=(A,b)\in\mathcal{P}$, where $A\in\mathbb{R}^{n\times n}$ is a non-singular matrix and $b\in\mathbb{R}^{n}$. The systems are derived from random generated matrices. Our code is simulated in Python and uses Pychop (Carson and Chen, 2025) for precision emulation.
We generated problems $Ax=b$ using two settings, each designed to evaluate the RL-driven precision selection for the GMRES-IR solver under different numerical conditions versus a full double precision routine (in the following, we use FP64 and double precision interchangeably). The agent with learning rate $\alpha$ of $0.5$ is trained within a custom Gym environment (Brockman et al., 2016) with 100 episodes, which simulates the process of solving a linear system using the GMRES-IR solver. To reduce the action space and accelerate computation, we adopted the reduction strategy outlined in Section 3.2 and further pruned it by selecting one-fourth of the valid precision combinations from the episode number combinations.
@@ 5.1 Problem Setup
where $|\mathcal{R}_{j}|$ is the cardinality $\mathcal{R}_{j}$, and the threshold $\tau_{j}$ is scaled by the condition number to account for problem difficulty, aligning with the error bounds in Demmel et al. (2006).
@@ 5.2 Simulations on Dense Systems
(a) RL ($W_{1}$, $\tau=10^{-6}$)
(b) RL ($W_{2}$, $\tau=10^{-8}$)
(c) RL ($W_{1}$, $\tau=10^{-6}$)
(d) RL ($W_{2}$, $\tau=10^{-8}$)
Table 5.2 reports error metrics and iteration counts for the linear solve, organized by condition number ranges: low ($10^{0}$ to $10^{3}$), medium ($10^{3}$ to $10^{6}$), and high ($10^{6}$ to $10^{9}$). Figure 2 details the precision types selected within finer condition number intervals.
@@
Regarding convergence, the conservative $W_{1}$ policy achieves universal convergence with a descent success rate $\xi=100\%$ across all settings. In comparison, the aggressive $W_{2}$ policy maintains $100\%$ success at $\tau=10^{-6}$ but shows reduced robustness at the stricter tolerance $\tau=10^{-8}$, with $\xi$ decreasing to $89.2\%$ in the low-$\kappa$ tests (RL fails in a few test samples). This is likely because the RL had insufficient exposure to training data in this specific range.
With respect to accuracy, $W_{1}$ closely matches the FP64 baseline, producing forward errors of the same order of magnitude and comparable backward errors. Precision lower than FP32 is never used in $W_{1}$. By contrast, $W_{2}$ exhibits a clear trade-off: BF16 and TF32 start to be used within low-to-medium $\kappa$ ranges, which is consistent with the fact that smaller condition numbers theoretically permit lower-precision computations with decreased error amplification. The $W_{2}$ policy permits the use of low precisions with substantially higher errors (e.g., ferr $\approx 10^{-7}$) to optimize the reward function, as long as the convergence quality is guaranteed.
This behavioral distinction also shows up in computational effort. The average number of outer iterations remains stable ($\approx 2.0$ to $2.35$) for both policies. However, the inner solver cost varies. The $W_{2}$ policy requires slightly more inner GMRES iterations (e.g., $8.00$ compared to $2.70$ for $W_{1}$ in low-$\kappa$ bins). This increase is needed to offset the growing rounding errors introduced by lower-precision formats.
The visualization of ferr and total GMRES iterations used for each test sample in $W_{2}$ is shown in Figure 3. The average performance degradation in Table 5.2 for the $W_{2}$ policy stems from a few instances that deviate from the identity line due to its aggressive strategy, favoring reduced precision.
The precision selection patterns in Figure 2 clarify these performance outcomes. In the low-$\kappa$ range ($10^{0}$ to $10^{3}$), $W_{1}$ employs a conservative combination of FP64 and FP32, entirely avoiding lower precisions. In contrast, $W_{2}$ leverages BF16 and TF32 (with usage frequencies of approximately $0.33$ and $0.76$, respectively), reflecting a focus on cost-efficiency. As the condition number increases, both strategies converge; beyond $10^{6}$, both policies adopt an FP64-dominant strategy and avoid the use of lower precisions.
@@ 5.3 Simulations on Sparse Systems
In this simulation, we emulated sparse linear systems following the experiment described in Häusner et al. (2024). A matrix $A_{0}\in\mathbb{R}^{n\times n}$ is constructed with $\text{nnz}(A_{0})=\lfloor\lambda_{s}n^{2}\rfloor$ non-zero entries, where $\lambda_{s}>0$ is the sparsity parameter. The entries are sampled from a standard normal distribution:
@@
The evaluation of sparse linear systems presents a distinct contrast to the dense case, primarily driven by the inherent numerical difficulty posed by the very ill-conditioned dataset. As depicted in Table 3, the test matrices exhibit high condition numbers between $1.02{\times}10^{8}$ and $7.86{\times}10^{9}$, which places linear solve in a range where numerical stability is precarious, leaving limited room for precision reduction. This test effectively demonstrates whether the RL agent can effectively avoid numerical instabilities in its precision selection.
Table 6: Average Performance Metrics Across Condition Ranges for Dense Systems (remove pernalty on iteration).
@@
5.4 Study on Iteration Penalty
The limitation of precision tuning in linear solvers without accounting for convergence iterations often comes with the most dynamic precision tuning tools like HiFPTuner (Guo and Rubio-González, 2018) because they focus only on precision tuning with the restriction on accuracy.
To further verify the effectiveness of our RL modeling for precision selection and performance balance, we further conduct a reward component analysis to evaluate the contribution of the penalty term. We focus on the total converged GMRES iterations taken for iterative refinement and conduct an ablation study on training with (21) without the penalty $f_{\text{penalty}}$ in the iteration terms.
(a) RL ($W_{1}$, $\tau=10^{-6}$)
(b) RL ($W_{2}$, $\tau=10^{-8}$)
(c) RL ($W_{1}$, $\tau=10^{-6}$)
(d) RL ($W_{2}$, $\tau=10^{-8}$)
Figure 4: Average Floating-point Types Selected Across Low, Medium, and High Condition Ranges (remove pernalty on iteration).
The results are presented in Table 5.3, together with Figure 4 showing precision types chosen by RL. Similarly to above, we observed that between low- and medium-conditioned linear systems, RL strategically selects aggressive reduced-precision steps, whereas in high-conditioned linear systems it adopts a more conservative strategy for precision selection. Compared to the training with $f_{\text{penalty}}$ as shown in Figure 2, the one without the $f_{\text{penalty}}$leads to more reduced-precision actions. The removal of it $f_{\text{penalty}}$ results in increasing iterations for the RL agent to achieve comparable accuracy, which means the iterative refinement uses extra iterations to compensate for the accuracy loss resulting from using increased low-precision steps. The results indicate the importance of the reward function modeling for RL training and demonstrate the effectiveness of RL for precision selection under proper reward modeling.